# Why Go Full? Elevating Federated Learning Through Partial Network Updates

### v2

@@ 3.4 Analysis for Communication and Computational Cost
Communication Cost. Suppose FedPart divides all layers into $M$ groups, and only one group is trained during each partial training session, with one communication round for each group. Then, the communication costs of FedPart will be $\frac{1}{M}$ of the original costs. This significant reduction can be explained as follows: Assume the communication cost of transmitting the complete model is $C$. By comparing the total communication cost for updating all parameters $Comm_{FNU}$ against the cost for updating only a subset of parameters $\boldsymbol{Comm_{PNU}}$ across one training cycle, we can derive:
$\frac{\boldsymbol{Comm_{PNU}}}{Comm_{FNU}}=\frac{C}{M*C}=\frac{1}{M}.$ (5)
Computational Cost. Assume computational costs are the same across all layers, our method in the partial network update phase can reduce the overall computational expense by $\frac{1}{3}$. The primary reason for this reduction is that in FedPart, it is unnecessary to compute gradients for the layers preceding the trainable parameters. To analyze this quantitatively, suppose the total overhead for both forward and backward propagation in a complete model satisfies $D_{tot}=D_{for}+D_{bak}$. The ratio of computational costs over a single training cycle can then be formulated as
@@
Given the nature of the backward propagation formulas, it is often approximated in the literature that the computational cost of backward propagation is about twice that of forward propagation (Rasley et al., 2020; Hobbhahn and Sevilla, 2021). Therefore, the above equation can be rewritten as:
@@ 4 Experiments
In the experimental setup, we primarily choose 40 clients, with local epochs to be 8. We test the global model on a balanced set. Unless specifically stated otherwise, the training datasets across all clients are independently and identically distributed (i.i.d.). We utilize the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001, which is determined to be the optimal learning rate. Consistent with prior references (Li et al., 2021b; Chen et al., 2022), we refrain from uploading local statistical information during model aggregation. Each experiment is conducted three times with different random seeds to ensure robustness. The experimental results for additional scenarios, including learning-rate tuning and client sampling, are presented in Appendix F.
When choosing experimental metrics, we employ three distinct measures to capture different facets of the benefits. These metrics include: Best Acc., which represents the ultimate accuracy achieved in classification tasks; Comm., indicating the total upstream transmission volume required by each client for a given training round (in GB); and Comp., which illustrates the total floating-point computation required by each client (in TFLOPs). All experiments are conducted on a server equipped with 8×A100 GPUs, and we provide the complete source code in our supplementary material.
@@ 4.1 Main Properties
The results in Table 1 show that our FedPart method demonstrates rapid convergence and consistently outperforms traditional FNU methods across all training cycles C, ultimately achieving significantly higher accuracy (e.g., improving FedAvg on Tiny-ImageNet by 21%). At the same time, its communication and computational costs are only 28% and 73% of those required by FNU. Furthermore, we observe that in some scenarios, the performance improvements of other federated learning algorithms even surpass those observed with FedAvg. This demonstrates that the layer mismatch problem we address is fundamentally different from issues tackled by previous works, and thus the performance enhancements they provide are independent of each other. This is a highly desirable property. However, our results on CIFAR-10 are less impressive. This suggests that in simpler datasets, the primary issue might be the client drift problem explored in previous studies, whereas the layer mismatch problem becomes more prominent in complex datasets.
@@
FedPart with deeper models. To evaluate the effectiveness of FedPart with deeper networks, we conduct experiments on ResNet-18 (detailed in Appendix A). This presents a more challenging scenario, as the proportion of trainable parameters significantly decreases in each round. Our experimental setup also follows the 2 R/L pattern, with five additional full network updates inserted between cycles. The results, displayed in Table 2, show that in deeper networks, FedPart not only maintains its advantages in terms of convergence speed and accuracy but also offers even more substantial reductions in communication and computational costs (by 85% and 27% compared to full network updates).
@@
FedPart for language modality. We also extend the FedPart method to the field of natural language processing and evaluate it on AGnews and SogouNews (Zhang et al., 2015) datasets. We choose the transformer architecture (Vaswani et al., 2017) for experiments. As shown in Table 3, the results indicate that FedPart performs well on language tasks, not only maintaining comparable performance as FNU, but also reducing communication and computational overhead by 66% and 25%, respectively. This demonstrates the method’s extensibility.
FedPart under data heterogeneity. We also evaluate the performance of FedPart under scenarios involving data heterogeneity. The results in Table 5 show that our FedPart consistently improve final performance(e.g., an improvement of 3.4% on Tiny-ImageNet) in the presence of data heterogeneity. However, the extent of performance improvement is relatively smaller. This suggests that client drift (Karimireddy et al., 2020) may have a more pronounced negative impact on our method. We also conduct experiments with extreme data heterogeneity ($\alpha=0.1$) in Appendix F.3.
@@ 4.2 Ablation Study
Training rounds per layer. In our FedPart, the training rounds per layer (denoted as $R/L$) is an important hyperparameter. A larger R/L value means more thorough training in each cycle, but it also results in a decrease in the number of cycles within the same number of training rounds. We explore the performance of FedPart under different R/L. From the results in Table 5, when $R/L=1$, the outcome shows limited final performance due to insufficient training for each layer. However, further increasing the R/L value not only fails to improve the final performance but also reduces the convergence speed. In extreme cases, when $R/L=10$, only one cycle is conducted overall, significantly impacting both the convergence speed and the final accuracy. This indicates that generally, increasing the number of cycles is more reasonable than extending their duration. This aligns with the motivation behind our proposal of multi-cycling training.
@@
Rounds of initial warm-up updates. To explore the impact of the duration of the initial full network updates phase (i.e. warm-up stage), we conduct experiments with this stage set to lengths of 0, 5, and 60. In Table 7, the term state refers to the period before or after partial network updates, which follow the warm-up phase. The experimental results clearly show that initial full network updates is crucial to the final model’s accuracy. However, extending the full network update phase yields diminishing returns. However, even when the model is trained with FNU until no further accuracy improvement is observed (60 init.), utilizing FedPart still enhances the model’s accuracy. This confirms FedPart’s capability to improve the convergence of the final global model and reduce layer mismatch.
Different orders for selecting trainable layers. We experiment with three different orders for selecting trainable parameters: sequential, reverse, and random. Sequential is the default configuration of FedPart, selecting layers from shallow to deep. In contrast, the reverse sequence selects layers from deep to shallow, while the random sequence selects layers randomly in each round. The results of the experiments are depicted in Table 7, demonstrating that the effectiveness of the three methods ranks as: sequential > reverse > random. This aligns with the intrinsic convergence order of neural networks and meets our experimental expectations.
@@ 4.3 Visualization Results
In this section, we conduct experiments to demonstrate why our proposed parameter selection strategy can enhance final performance, and what the impact of layer-wise information exchange has on privacy. Our experiments are based on ResNet-8 and the CIFAR-100 dataset. We analyze the models obtained from four different methods: 1) FedAvg-100, which represents training with full network for 100 rounds; 2) FedPart(No Init. 1C), which represents using FedPart for one cycle without initial full network updates; 3) FedPart(1C), which involves initial full network updates followed by one cycle of FedPart training; 4) FedPart(5C), which involves initial full network updates followed by five cycles of FedPart training. The visualization results are as follows.
@@
Convolutional kernel visualization. We also analyze how different models extract semantic information by visualizing the convolutional kernels. We find that in the full network updates represented by FedAvg-100, the shallow convolutional kernels primarily function as edge/corner detectors. However, direct training of partial networks disrupts this property. Further, by employing initial full network updates and adding multiple training cycles, we gradually restore this characteristic. This effectively explains the impact of the parameter selection strategy on the final model formation. For specific visualization results of the convolutional kernels, please refer to Appendix D.
4.4 Privacy Issue
We further posit that FedPart exhibit stronger ability of privacy protection, as it transmits less information in each communication round. Formally, we can abstract the model training process (for both full and partial parameter training) as a mapping: $(\Delta w_{1},\Delta w_{2},...,\Delta w_{n})=f(x)$, where the left hand side denotes the updates to each model parameter, and $x$ is the training data. From a privacy attack perspective, the goal is to find the best $x$ such that the updates to $w$ are as close as possible to the actual updates in each dimension. This resembles solving a system of equations, where $x$ are the unknowns, and each dimension of $w$ update represents an equation. With partial network training, the unknowns $x$ remain unchanged compared with full parameter training, but the number of equations decreases (i.e., less information for the attack to follow). Therefore, we believe partial network training leakages less information in general.
@@ 4.4 Privacy Issue
And we use PSNR to measure the quality of the reconstructed image. Since $x$ denote the original image and $\hat{x}$ the reconstructed image, the PSNR is calculated as follows:
@@
A smaller MSE value indicates a higher similarity between the two images. Consequently, a larger PSNR value implies a better quality of the reconstructed image.
The results in Table 9 show that, for different trainable layers, our method consistently exhibits better privacy protection in both average and worst-case scenarios compared to full network updates. Attacking examples and detailed metric implementations are provided in Appendix E.
@@ 5 Conclusion and Limitation
This article identifies that the model averaged in federated learning is not directly applicable to the specific tasks of each client, a situation we refer to as layer mismatch. To address this issue, we propose the FedPart method, which introduces a strategy for selecting and training partial networks. We validate the effectiveness of FedPart both theoretically and experimentally. In future work, we plan to evaluate our method on a wider range of model architectures and apply it to larger-scale datasets to further investigate its effectiveness and scalability.
6 Acknowledgement
@@ Appendix B Proof for Convergence Rate of FedPart
According to Assumption 2, the upper limit of the left side of the inequality is $\sigma^{2}$, so we finally obtain a general upper bound for the gradient with masks:
@@
Having prepared the groundwork, we are now ready to begin the formal proof process. First, based on Assumption 1, as the loss function is L-smooth, we have:
@@
Further expanding the right-hand side of the above inequality, we can obtain:
@@
In the above derivation, we have used the assumption of L-smoothness and Jensen’s inequality. Next, we will continue to estimate the upper limit of this term. Assuming that the last time $t=t_{0}$ was a parameter aggregation, and the next time the parameters aggregate is at $t=t_{0}+E$, then:
@@
Finally, summing the inequalities from $t=1,\cdots,T$, and multiplying both sides by $\frac{1}{T}$, we can obtain:
@@
Selecting a learning rate $\gamma=\frac{\sqrt{NM}}{L\sqrt{T}}$, we can obtain: $\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[||S_{i}^{t}\odot\nabla f(\bar{w}^{t-1})||^{2}\leq\frac{2L}{\sqrt{NMT}}(f(\bar{w}^{0})-f^{*})+\frac{4NME^{2}G^{2}}{T}+\frac{\sigma^{2}k^{2}}{\sqrt{NMT}}$. Further choosing $E\leq\frac{T^{1/4}}{(MN)^{3/4}}$, we can obtain a further corollary: $\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[||S_{i}^{t}\odot\nabla f(\bar{w}^{t-1})||^{2}\leq\frac{2L}{\sqrt{NMT}}(f(\bar{w}^{0})-f^{*})+\frac{4G^{2}}{\sqrt{MNT}}+\frac{\sigma^{2}k^{2}}{\sqrt{NMT}}=O(\frac{1}{\sqrt{NMT}})$. This proves the convergence rate of FedPart.
@@ Appendix C Visualizations for Activation Maximization
From the visualization results, it can be observed that FedAvg-100, due to being a full network update, captures low-level semantic features (such as clear boundaries) in shallow layers, while deeper layers capture complex semantic information. However, the results of FedPart(No Init. 1C) exhibit noticeable differences in color and structural features compared to the full network update. This verifies our conjunction that partial network update is detrimental to forming such a hierarchical information extraction approach, leading to the model converging to possible local minima. Additionally, we can also see that by including the initial phase of full network updates and multiple rounds of sequential training, the similarity of semantic information obtained by the model gradually approaches that of FedAvg. Therefore, the results sufficiently demonstrate that although we only train one layer of the network each time, by employing an appropriate layer selection scheme, we ultimately achieve an effect close to that of full network updates.
@@ Appendix E Robustness to Privacy Attack
In this section, we will provide a detailed introduction to the results of privacy leakage using the DLG method on full network updates and partial network updates. We perform DLG attacks in four settings: transmiting all parameters in FedAvg-100 model; and transmiting only the parameters of layers #$\#1$, #$\#9$, and #$\#10$ separately in FedPart(5C) model. In Fig. 8, we select some representative reconstructed images. The leftmost column represents the original images, while the four columns on the right show the reconstructed images obtained through DLG attacks under different settings.

### v3

@@ 3.4 Analysis for Communication and Computational Cost
Communication Cost. Suppose FedPart divides all layers into $M$ groups, and only one group is trained during each partial training session, with one communication round for each group. It is easy to verify that the averaged communication costs of FedPart during the partial network update phase is reduced to $\frac{1}{M}$ of the original costs.
Computational Cost. Assuming that computational costs are uniform across all layers, our method can reduce the overall computational expense during the partial network update phase by $\frac{1}{3}$. The primary reason for this reduction is that in FedPart, there is no need to compute gradients for the layers preceding the trainable parameters. To analyze this quantitatively, suppose the total overhead for both forward and backward propagation in a complete model satisfies $D_{tot}=D_{for}+D_{bak}$. The ratio of computational costs over a single training cycle can then be written as
@@
Moreover, it is widely accepted in the literature that the computational cost of backward propagation is approximately twice that of forward propagation (Rasley et al., 2020; Hobbhahn and Sevilla, 2021). Therefore, the above equation can be rewritten as:
@@ 4 Experiments
In the experimental setup, we primarily choose 40 clients, with local epochs to be 8. We test the global model on a balanced set. Unless specifically stated otherwise, the training datasets across all clients are independently and identically distributed (i.i.d.). We utilize the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001, which is determined to be the optimal learning rate. In line with prior references (Li et al., 2021b; Chen et al., 2022), we refrain from uploading local statistical information during model aggregation. Each experiment is conducted three times with different random seeds to ensure robustness. The experimental results for additional scenarios, including learning-rate tuning and client sampling, are presented in Appendix F.
When choosing experimental metrics, we employ three distinct measures to capture various aspects of the benefits. These metrics include: Best Acc., which represents the ultimate accuracy achieved in classification tasks; Comm., indicating the total upstream transmission volume required by each client for a given training round (in GB); and Comp., which illustrates the total floating-point computation required by each client (in TFLOPs). All experiments are conducted on a server equipped with 8×A100 GPUs, and we provide the complete source code in our supplementary material.
@@ 4.1 Main Properties
The results in Table 1 show that our FedPart method demonstrates rapid convergence and consistently outperforms traditional FNU methods across all training cycles C, ultimately achieving significantly higher accuracy (e.g., improving FedAvg on Tiny-ImageNet by 21%). At the same time, its communication and computational costs are only 28% and 73% of those required by FNU. Furthermore, we observe that in some scenarios, the performance improvements of other federated learning algorithms even surpass those observed with FedAvg. This highlights that the layer mismatch problem identified in this paper is novel and cannot be addressed by any existing methods. However, our results on CIFAR-10 are less impressive. This suggests that in simpler datasets, the primary issue might be the client drift problem explored in previous studies, whereas the layer mismatch problem becomes more prominent in complex datasets.
@@
FedPart with deeper models. To evaluate the effectiveness of FedPart with deeper networks, we conduct experiments on ResNet-18 (detailed in Appendix A). This presents a more challenging scenario, as the proportion of trainable parameters significantly decreases in each round. Our experimental setup also follows the 2 R/L pattern, with five additional full network updates inserted between cycles. The results, displayed in Table 2, show that in deeper networks, FedPart not only maintains its advantages in convergence speed and accuracy but also provides even greater reductions in communication and computational costs (by 85% and 27% compared to full network updates).
@@
FedPart for language modality. We also extend the FedPart method to the field of natural language processing and evaluate it on AGnews and SogouNews (Zhang et al., 2015) datasets. We choose the transformer architecture (Vaswani et al., 2017) for experiments. As shown in Table 3, the results indicate that FedPart performs well on language tasks, not only maintaining comparable performance as FNU, but also reducing communication and computational overhead by 66% and 25%, respectively. This demonstrates the method’s scalability.
FedPart under data heterogeneity. We also evaluate the performance of FedPart under scenarios involving data heterogeneity. The results in Table 5 show that our FedPart consistently improve final performance (e.g., an improvement of 3.4% on Tiny-ImageNet) in the presence of data heterogeneity. However, the extent of performance improvement is relatively smaller. This suggests that client drift (Karimireddy et al., 2020) may have a more pronounced negative impact on our method. We conduct experiments with extreme data heterogeneity ($\alpha=0.1$) in Appendix F.3.
@@ 4.2 Ablation Study
Training rounds per layer. In our FedPart, the training rounds per layer (denoted as R/L) is an important hyperparameter. A larger R/L value means more thorough training in each cycle, but it also results in a decrease in the number of cycles within the same number of training rounds. We explore the performance of FedPart under different R/L. From the results in Table 5, when R/L=1, the outcome shows limited final performance due to insufficient training for each layer. However, further increasing the R/L value not only fails to improve the final performance but also reduces the convergence speed. In extreme cases, when R/L=10, only one cycle is conducted overall, which significantly affects both convergence speed and final accuracy. This indicates that generally, increasing the number of cycles is more effective than extending their duration. This aligns with the motivation behind our proposal of multi-cycling training.
@@
Rounds of initial warm-up updates. To explore the impact of the duration of the initial full network updates phase (i.e. warm-up stage), we conduct experiments with this stage set to lengths of 0, 5, and 60. In Table 7, the term state refers to the period before or after partial network updates, which follow the warm-up phase. The experimental results clearly show that initial full network updates is crucial to the final model’s accuracy. Notably, extending the full network update phase yields diminishing returns. However, even when the model is trained with FNU until no further accuracy improvement is observed (60 init.), utilizing FedPart still enhances the model’s accuracy. This confirms FedPart’s capability to improve the convergence of the final global model and reduce layer mismatch.
Different orders for selecting trainable layers. We experiment with three different orders for selecting trainable parameters: sequential, reverse, and random. Sequential is the default configuration of FedPart, selecting layers from shallow to deep. In contrast, the reverse sequence selects layers from deep to shallow, while the random sequence selects layers randomly in each round. The results of the experiments are depicted in Table 7, demonstrating that the effectiveness of the three methods ranks as follows: sequential > reverse > random. This aligns with the intrinsic convergence order of neural networks and meets our experimental expectations.
@@ 4.3 Visualization Results
In this section, we conduct experiments to demonstrate why our proposed parameter selection strategy can enhance final performance, and what the impact of layer-wise information exchange has on privacy leakage. Our experiments are based on ResNet-8 and the CIFAR-100 dataset. We analyze the models obtained from four different methods: 1) FedAvg-100, which represents training with full network for 100 rounds; 2) FedPart(No Init. 1C), which represents using FedPart for one cycle without initial full network updates; 3) FedPart(1C), which involves initial full network updates followed by one cycle of FedPart training; 4) FedPart(5C), which involves initial full network updates followed by five cycles of FedPart training. The visualization results are as follows.
@@
Convolutional kernel visualization. We also analyze how different models extract semantic information by visualizing the convolutional kernels. We find that in the full network updates represented by FedAvg-100, the shallow convolutional kernels primarily function as edge/corner detectors. However, direct training of partial networks disrupts this property. Further, by initially employing full network updates and adding multiple training cycles, we gradually restore this characteristic. This effectively explains the impact of the parameter selection strategy on the final model formation. For specific visualization results of the convolutional kernels, please refer to Appendix D.
4.4 Impact on Privacy Protection
We next demonstrate that FedPart offers enhanced privacy protection, as it transmits less information in each communication round. Formally, we can abstract the model training process (for both full and partial parameter training) as a mapping: $(\Delta w_{1},\Delta w_{2},...,\Delta w_{n})=f(x)$, where the left hand side denotes the updates to each model parameter, and $x$ is the training data. From a privacy attack perspective, the goal is to find the best $x$ such that the updates to $w$ are as close as possible to the actual updates in each dimension. This resembles solving a system of equations, where $x$ are the unknowns, and each dimension of $w$ update represents an equation. With partial network training, the unknowns $x$ remain unchanged compared with full parameter training, but the number of equations decreases (i.e., less information for the attacker to leverage). Therefore, we believe that partial network training generally leaks less information.
@@ 4.4 Impact on Privacy Protection
We use PSNR to measure the quality of the reconstructed image. Given that $x$ denotes the original image and $\hat{x}$ denotes the reconstructed image, the PSNR is calculated as follows:
@@
A larger PSNR value means a better quality of the reconstructed image, which further implies a higher risk of privacy leakage.
The results in Table 9 show that, for different trainable layers, our method consistently exhibits better privacy protection in both average and worst-case scenarios compared to full network updates. Attacking examples are provided in Appendix E.
@@ 5 Conclusion and Limitation
We observe that the model averaged in federated learning is not directly applicable to the specific tasks of each client, a situation we refer to as layer mismatch. To address this issue, we propose the FedPart method, which introduces a strategy for selecting and training partial networks. We validate the effectiveness of FedPart both theoretically and experimentally. In future work, we plan to evaluate our method on a wider range of model architectures and apply it to larger-scale datasets to further investigate its effectiveness and scalability.
Acknowledgement
@@ Appendix B Proof for Convergence Rate of FedPart
For following proof, we mainly refer to Yu et al. [2019]. According to Assumption 2, the upper limit of the left side of the inequality is $\sigma^{2}$, so we finally obtain a general upper bound for the gradient with masks:
@@
With the groundwork laid, we are now ready to begin the formal proof process. First, based on Assumption 1, as the loss function is L-smooth, we have:
@@
Further expanding the right-hand side of the above inequality, we obtain:
@@
In the above derivation, we have used the assumption of L-smoothness and Jensen’s inequality. Next, we will continue to estimate the upper limit of this term. Assuming that the last parameter aggregation occurred at time $t=t_{0}$, and the next aggregation will take place at $t=t_{0}+E$, then:
@@
Finally, summing the inequalities from $t=1,\cdots,T$, and multiplying both sides by $\frac{1}{T}$, we obtain:
@@
Selecting a learning rate $\gamma=\frac{\sqrt{NM}}{L\sqrt{T}}$, we obtain: $\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[||S_{i}^{t}\odot\nabla f(\bar{w}^{t-1})||^{2}\leq\frac{2L}{\sqrt{NMT}}(f(\bar{w}^{0})-f^{*})+\frac{4NME^{2}G^{2}}{T}+\frac{\sigma^{2}k^{2}}{\sqrt{NMT}}$. Furthermore, by choosing $E\leq\frac{T^{1/4}}{(MN)^{3/4}}$, we can derive the following corollary: $\frac{1}{T}\sum_{i=1}^{T}\mathbb{E}[||S_{i}^{t}\odot\nabla f(\bar{w}^{t-1})||^{2}\leq\frac{2L}{\sqrt{NMT}}(f(\bar{w}^{0})-f^{*})+\frac{4G^{2}}{\sqrt{MNT}}+\frac{\sigma^{2}k^{2}}{\sqrt{NMT}}=O(\frac{1}{\sqrt{NMT}})$. This proves the convergence rate of FedPart.
@@ Appendix C Visualizations for Activation Maximization
From the visualization results, it can be observed that FedAvg-100, due to being a full network update, captures low-level semantic features (such as clear boundaries) in shallow layers, while deeper layers capture complex semantic information. However, the results of FedPart(No Init. 1C) exhibit noticeable differences in color and structural features compared to the full network update. This confirms our belief that partial network updates are detrimental to establishing a hierarchical information extraction approach, resulting in the model converging to possible local minima. Additionally, we observe that by including the initial phase of full network updates and multiple rounds of sequential training, the similarity of semantic information obtained by the model gradually approaches that of FedAvg. Therefore, the results sufficiently demonstrate that although we only train one layer of the network each time, by employing an appropriate layer selection scheme, we ultimately achieve results comparable to those of full network updates.
@@ Appendix E Robustness to Privacy Attack
In this section, we will examine the privacy leakage using the DLG method on full network updates and partial network updates. We perform DLG attacks in four settings: transmiting all parameters in FedAvg-100 model; and transmiting only the parameters of layers #$\#1$, #$\#9$, and #$\#10$ separately in FedPart(5C) model. In Fig. 8, we select some representative reconstructed images. The leftmost column represents the original images, while the four columns on the right show the reconstructed images obtained through DLG attacks under different settings.